OpenStudy (nicoleced):

9a^2 * 5/36 <-fraction

OpenStudy (kl0723):

is it: \[9a^2*\frac{ 5 }{ 36 }?\]

OpenStudy (nicoleced):

yes

OpenStudy (kl0723):

ok, what is your question?

OpenStudy (nicoleced):

how to calculate it

OpenStudy (nicoleced):

i mean solve the problem *

OpenStudy (kl0723):

are you trying to reduce to a simple term or solving for a?

OpenStudy (nicoleced):

multiply and simplify

OpenStudy (nicoleced):

@kl0723

OpenStudy (kl0723):

there you go :) it shouldn't be that difficult

OpenStudy (nicoleced):

how did you get 5 for the numerator?

OpenStudy (nicoleced):

@kl0723

OpenStudy (kl0723):

@nicoleced ohh ok, notice that the numerator and the denominator are divisible by the number 9 (sorry I wrote 4) :( so you have simplify by extracting 9 from top and bottom

OpenStudy (kl0723):

you now will have, \[\frac{ 45 = (5*9) }{ 36=(4*9) } \] in other words you simplify that 9 from both top and bottom, leaving you with the remaining \[\frac{ 5 }{ 4 }\]

OpenStudy (kl0723):

the 9 on the top will cancel out with the 9 on the bottom :)

OpenStudy (nicoleced):

yess, i get it now! thanks! :)